# Unbiased General Annotated Dataset Generation

### v1

@@ Abstract
Pre-training backbone networks on a general annotated dataset (e.g., ImageNet) that comprises numerous manually collected images with category annotations has proven to be indispensable for enhancing the generalization capacity of downstream visual tasks. However, those manually collected images often exhibit bias, which is non-transferable across either categories or domains, thus causing the model’s generalization capacity degeneration. To mitigate this problem, we present an unbiased general annotated dataset generation framework (ubGen). Instead of expensive manual collection, we aim at directly generating unbiased images with category annotations. To achieve this goal, we propose to leverage the advantage of a multimodal foundation model (e.g., CLIP), in terms of aligning images in an unbiased semantic space defined by language. Specifically, we develop a bi-level semantic alignment loss, which not only forces all generated images to be consistent with the semantic distribution of all categories belonging to the target dataset in an adversarial learning manner, but also requires each generated image to match the semantic description of its category name. In addition, we further cast an existing image quality scoring model into a quality assurance loss to preserve the quality of the generated image. By leveraging these two loss functions, we can obtain an unbiased image generation model by simply fine-tuning a pre-trained diffusion model using only all category names in the target dataset as input. Experimental results confirm that, compared with the manually labeled dataset or other synthetic datasets, the utilization of our generated unbiased datasets leads to stable generalization capacity enhancement of different backbone networks across various tasks, especially in tasks where the manually labeled samples are scarce. Code: https://github.com/vvvvvjdy/ubGen
@@ 1 Introduction
Deep neural networks have achieved great success in various computer vision tasks [34, 61, 37]. One indispensable premise of such success lies on pre-training the parameter-extensive backbone network using a general annotated dataset (e.g., ImageNet [8]) that contains a large number of images with manually annotated categories. Profiting from the vast amount of annotated images and the diverse image categories, the pre-trained backbone networks often show pleasing generalization capacity and perform effectively in the target computer vision task through simple fine-tuning it together with a corresponding task head with a few parameters [17, 57]. Unfortunately, recent studies [30, 51] uncover that these manually collected images often exhibit non-travail bias (e.g., a certain background, image style, object position for a specific category etc.) which can be easily captured by backbone networks during pre-training, but hardly noticed by human collectors (see Fig. 1). Such hidden bias is pone to be cast into a shortcut feature representation [14] to improve the in-domain performance but deteriorate the generalization capacity of pre-trained backbone networks on target tasks in the cross-category or cross-domain settings [12, 4] which shows an obvious image distribution gap from the utilized general annotated dataset. For example, when a specific category of images often shows a similar background, the pre-train backbone networks will consider the background as the discriminative feature of such a category while overlooking the cross-category or cross-domain transferable semantic features (e.g., shape, structure, etc.). Therefore, it is crucial to obtain an unbiased general annotated dataset to enhance the cross-category or cross-domain generalization capacity of the pre-trained backbone networks.
@@
To this end, a straight solution is to manually re-collect extensive unbiased images. However, it will not only produce expensive manual collection costs, but also inevitably incur some other undetectable bias. Recently, diffusion models [44, 43, 41] have shown powerful capacity in terms of generating high-quality synthetic images based on the text description of image contents, thus providing a feasible way to directly generate images with annotations without manual collection cost. Moreover, some studies [48, 60, 1, 46] have demonstrated that those randomly generated images with annotations can be utilized for network training. Although most existing diffusion models can be directly utilized for general annotated dataset generation, they mainly focus on generating images with the distribution consistent with the conventional manually annotated general dataset (e.g., ImageNet) and scarcely attempt to generate unbiased images. Thus, pre-training the backbone networks on these generated general annotated datasets will not bring non-travail generalization capacity enhancement.
@@
To mitigate this problem, we present an unbiased general annotated dataset generation framework (ubGen), which takes the first attempt to directly generating synthetic unbiased images with category annotations. To achieve this goal, we first have to define an unbiased space where the feature representation of each image emphasizes the cross-category or cross-domain transferable semantic characteristics. Considering the recent progress of multimodal foundation model (e.g., CLIP [40]) which aims at mapping images into an unbiased semantic space defined by language, a straight idea is to constrain the image output of the existing diffusion models to follow the semantic distribution of the specific image category in such an unbiased semantic space. Following this idea, we develop a bi-level semantic alignment loss based on the CLIP model to fine-tune the pre-trained diffusion model. In a specific, on one hand, such a loss forces all generated images to be consistent with the semantic distribution of all categories belonging to the target dataset in the CLIP feature space using an adversarial learning scheme. On the other hand, it also requires each generated image to match the semantic description of its category name in the CLIP feature space using a simple cosine similarity metric. By doing these, we can obtain an unbiased image generation model by simply fine-tuning a pre-trained diffusion model using only all category names in the target dataset as input. In addition, to sidestep image quality degradation caused by the unbiased image generation learning, we further cast an existing image quality scoring model into a quality assurance loss to assist the bi-level semantic alignment loss for diffusion model fine-tuning.
@@
To testify the efficacy of the proposed framework, we pre-train two conventional backbone networks on our generated unbiased general annotated dataset and training specific heads on different downstream tasks. Compared with backbone networks pre-trained either on the manually collected generated dataset or that generated by existing diffusion models, our approach achieves obvious generalization performance improvement, especially when the manually annotated samples in the target task are scarce. Moreover, additional experiments prove that our pre-trained backbone networks capture lower specific bias (e.g. context, background, shape-texture), which further demonstrates the generality of our framework.
@@
  • We propose the first unbiased general annotated dataset generation framework, which jumps out of the dilemma of traditional manual data collection in terms of mitigating dataset bias.
@@
  • With our generated general unbiased dataset, the pre-trained backbone network shows state-of-the-art generalization capacity in different downstream tasks.
@@ 2.1 Datasets Bias
Since the deep learning revolution in 2012 [24], the large-scale manually collected annotated dataset (e.g., ImageNet) performs no longer a simple training dataset for its own tasks, but a general dataset utilized for backbone network pre-training in the universal computer vision world. Moreover, pre-training backbone networks on a general annotated dataset has become the indelible step for enhancing the generalization performance of various downstream tasks in practice. However, recent studies have consecutively revealed that these existing manually collected general dataset exhibit non-trial bias, which results in sup-optimal cross-categories and cross-domain generalization capacity, especially when the manually annotated samples in the target task are scarce. For example, Liu et al. [30] observe that deep neural networks can achieve excellent accuracy in classifying which dataset an image is from. In other words, the neural network discovers some dataset-specific patterns, a form of bias. In addition, studies in [48, 19, 37] attempt to implicitly measure the dataset bias by investigating the cross-category or cross-domain generalization capacity as well as the robustness of the models pre-trained on the dataset. These works mainly focus on raising the dataset bias problem or bias measurement. In contrast, in this study, we take the first attempt to solve this problem and aims at borrowing the advantage of diffusion model in image generation to directly generate a unbiased general dataset for better backbone pre-training.
@@ 2.2 Synthetic Dataset Generation
Different from manual collection, synthetic dataset generation aims at directly generating image using deep neural networks based on some text description without any manual annotation cost. For example, in early days, Zhu et al. [63, 58] utilized adversarial generative networks to model the mapping relationship between the input text description and the output image and is able to randomly generate various images. However, these methods require large-scale high-quality images from target categories for network training, and the generalization capacity to unknown text descriptions is limited. To mitigate this problem, recent studies mainly focus on the large-scale diffusion model which is pre-trained on a great number of images and shows powerful capacity in generalization to unknown text description [43, 38, 11]. Thus, some recent works have attempted to utilize the diffusion model to generate ImageNet-like synthetic dataset for backbone pre-training. For example, Bansal et al. [1] directly fine-tune the diffusion model on ImagNet-1K [8] and use meticulously designed prompts to generate the dataset. Lei et al. [26] utilizes ViT-GPT2 [25] to get a unique prompt to generate each image. Yuan et al. [60] resort to learn the distribution of ImageNet and use Blip-captions [27] of ImageNet as prompts to synthesize the dataset. Although these diffusion-based methods can be utilized for general dataset generation, they mainly focus on simulating the existing ImageNet without considering the dataset bias. In this study, we attempt to fine-tune the diffusion model to directly generate an unbiased general annotated dataset without using any image-text pairs but the category names of the target dataset as input.
@@ 3 Approach
The overall training framework of ubGen is shown in Fig. 2. In Sec. 3.1, we begin with elucidating the basic methodology employed in our training process. In Sec. 3.2, we then illustrate the bi-level semantic alignment module, which is the core part of our approach. Subsequently, in Sec. 8, we further introduce the quality assurance module for fidelity preserving the images, and finally we integrate the two components for joint learning.
@@ 3.2 Bi-Level Semantic Alignment
As mentioned in Sec. 1, we assume that the semantic space defined by language can be a unbiased representation and our key insight is using this characteristic of language to fine-tune a pre-trained diffusion model as our ubGen generator. We achieve this by leveraging a simple Linear-ReLU-Linear based discriminator $\mathcal{D}_{\phi}$ and utilizing CLIP to carry out a bi-level semantic alignment. We use only 1000 class names of ImageNet as our inputs, which ensures that no other biased information is introduced except semantic information of the dataset.
Entire Dataset Alignment. Building on the advantages of CLIP which has unified the image and the text into one represention space, we can initially use CLIP text encoder to extract text features $\{f_{c_{1}},f_{c_{2}},\dots,f_{c_{1000}}\}$ of classnames $\{c_{1},c_{2},\dots,c_{1000}\}$. We consider these text features to be an unbiased semantic distribution of the entire ImageNet. Then, we generate an image using prompt $c_{i}$ and send it to CLIP image encoder to get the image feature $f_{im_{i}}$. Next, we randomly choose a text feature $f_{c_{j}}$ from $\{f_{c_{1}},f_{c_{2}},\dots,f_{c_{1000}}\}$. It is worth emphasizing that we do not use the text feature which belongs to the same class as the image feature since we aim to align the whole synthetic dataset to its general semantic representation space regardless of the concrete class. Finally, features $f_{im_{i}}$ and $f_{c_{j}}$ are fed into $\mathcal{D}_{\phi}$ for computing entire semantic alignment loss as follows:
@@
Individual Image Alignment. Expect for mapping the images to be consistent with the semantic distribution of all classes within the entire dataset, we need to precisely control each category of images to match their semantic description. To this end, we introduce the individual semantic alignment loss. In particular, given the generated image ${im_{i}}$ using class name $c_{i}$, we use simple "photo of $c_{i}$" as the unbiased semantic description $p_{c_{i}}$ and send them to CLIP. Different from using CLIP to align the dataset globally in the entire semantic space, we aim at forcing the semantic information of each image to be dovetailed with its class by maximizing the cosine similarity between the image and its corresponding semantic description through CLIP. Thus, we can obtain $\mathcal{L}_{in}$ formulated as follows:
@@
By considering these two distinct levels of losses, we can finally add them together and obtain bi-level semantic alignment loss $\mathcal{L}_{bi}$ to refine the diffusion model to align more closely with the unbiased semantic reference.
@@ 3.3 Quality Assurance
Finally, we combine the losses in the bi-level semantic alignment module and quality assurance module to build up our final training object for ubGen generator as follows:
$\mathcal{L}=\mathcal{L}_{bi}+\lambda_{1}\mathcal{L}_{q}$ (8)
@@
IN-ubGen(ours) $46.3$ $62.6$ $58.0$ $71.2$ $94.1$ $86.2$ $88.6$ $68.5$ $66.0$ $74.4$
Table 1: Top-1 accuracy on transfer learning datasets. We utilize the pre-trained visual backbones as feature extractors and train linear classifiers to test their generalization ability. The average accuracy across eight transfer learning datasets is denoted as Avg. The best and second-best transfer learning performance of each backbone are highlighted in red and underlined. IN-SD1.5 denotes only using original SD1.5 to generate the data. We also present results on ImageNet validation set for reference.
@@ 4.1 Experimental Settings
More details about training hypersettings, data synthesis, and datasets are provided in Appendix.
@@ 4.2 Generalization across Downstream Tasks
Transfer Learning. Transfer learning [37] is a widely known downstream visual task and can be significantly influenced by the generalization of the pre-trained model. In our work, we aim to indicate whether the utilization of our unGen data can enable the backbones to learn better cross-category transferable patterns. To this end, we follow fakeit [46], which uses pre-trained visual backbones as feature extractors and train simple linear logistic regression classifiers from scratch.
Figure 3: Scaling down the number of training images of eight transfer learning datasets. The benefits of using pre-trained models on our ubGen images are even more pronounced when there is less data for training.
@@
As illustrated in Tab. 1, we observe that models pre-trained on our dataset outperform all candidates. Compared with the second-best synthetic datasets, we achieve $+1.4\%$ and $+3.6\%$ leading performance on ResNet50 and ViT-S respectively. More importantly, our ubGen dataset exhibits $1.7\%$ and $2.1\%$ average accuracy improvement and superior results in the vast majority of transfer learning datasets compared to real data. Furthermore, we empirically investigate the transfer ability of models pre-trained on real images and our generated images when using less transfer learning training data. Such few-shot setting [52] requires an even higher generalization capacity of the model. As shown in Fig.3, it is important to note that the advantage of pre-trained models on our ubGen-ImageNet is even greater when there are fewer downstream images for training. This phenomenon further underscores the lower bias of our data and the stronger generalization of the resulting model. Meanwhile, another striking finding is that achieving high accuracy on the ImageNet validation set does not necessarily correlate with enhanced cross-category generalization performance, thereby enabling us to draw more definitive conclusions regarding the dualistic impact of bias in the existing dataset on the model.
Visual Perception. Detection and segmentation are two of the most popular downstream visual perception tasks, at the same time, these two tasks can benefit a lot from a well pre-trained backbone. Hence, We want to find out if the utilization of our ubGen data can also enhance the performance on these tasks. To this end, we follow previous studies [31, 32, 47], which use Mask R-CNN [17] as detection head for COCO object detection and UperNet [57] as segmentation head for ADE20K semantic segmentation, to evaluate the performance of pre-trained ResNet50 backbone (ViT do not support higher resolution images) on different datasets. Moreover, to thoroughly test the generalization ability of the backbone, we progressively decrease the number of training samples and observe the outcomes.
The experimental results in Table 2 demonstrate the effectiveness of our ubGen datasets. Although pre-training on real data achieves marginally better performance with full training data, the performance of ubGen pre-trained model consistently outperforms all other pre-training data when downstream data is limited. Specifically, with only $20\%$ of the original training data, the model achieves the highest performance across both tasks, showing gains of up to $1.54\%$ APbox and $1.47\%$ mIoU compared to that pre-trained on IN-Real. This result is particularly valuable for real-world applications where collecting and annotating task-specific training data is often costly and time-consuming.

### v3

# Low-Biased General Annotated Dataset Generation

@@ Abstract
Pre-training backbone networks on a general annotated dataset (e.g., ImageNet) that comprises numerous manually collected images with category annotations has proven to be indispensable for enhancing the generalization capacity of downstream visual tasks. However, those manually collected images often exhibit bias, which is non-transferable across either categories or domains, thus causing the model’s generalization capacity degeneration. To mitigate this problem, we present a low-biased general annotated dataset generation framework (lbGen). Instead of expensive manual collection, we aim at directly generating low-biased images with category annotations. To achieve this goal, we propose to leverage the advantage of a multimodal foundation model (e.g., CLIP), in terms of aligning images in a low-biased semantic space defined by language. Specifically, we develop a bi-level semantic alignment loss, which not only forces all generated images to be consistent with the semantic distribution of all categories belonging to the target dataset in an adversarial learning manner, but also requires each generated image to match the semantic description of its category name. In addition, we further cast an existing image quality scoring model into a quality assurance loss to preserve the quality of the generated image. By leveraging these two loss functions, we can obtain a low-biased image generation model by simply fine-tuning a pre-trained diffusion model using only all category names in the target dataset as input. Experimental results confirm that, compared with the manually labeled dataset or other synthetic datasets, the utilization of our generated low-biased dataset leads to stable generalization capacity enhancement of different backbone networks across various tasks, especially in tasks where the manually labeled samples are scarce. Code is available at: https://github.com/vvvvvjdy/lbGen
@@ 1 Introduction
Deep neural networks have achieved great success in various computer vision tasks [34, 62, 37]. One indispensable premise of such success lies on pre-training the parameter-extensive backbone network using a general annotated dataset (e.g., ImageNet [8]) that contains a large number of images with manually annotated categories. Profiting from the vast amount of annotated images and the diverse image categories, the pre-trained backbone networks often show pleasing generalization capacity and perform effectively in the target computer vision task through simple fine-tuning it together with a corresponding task head with a few parameters [17, 57]. Unfortunately, recent studies [30, 51, 61] uncover that these manually collected images often exhibit non-travail bias [^1] In this paper, ‘dataset bias’ refers to ‘systematic bias introduced in data collection, selection, or processing that impair the generalization capacity of the model’. (e.g., a certain background, image style, object position for a specific category etc.) which can be easily captured by backbone networks during pre-training, but hardly noticed by human collectors (see Figure 1). Such hidden bias is proven to be cast into a shortcut feature representation [14] to improve the in-domain performance but deteriorate the generalization capacity of pre-trained backbone networks on target tasks in the cross-category or cross-domain settings [12, 4] which shows an obvious image distribution gap from the utilized general annotated dataset. For example, when a specific category of images often shows a similar background, the pre-train backbone networks will consider the background as the discriminative feature of such a category while overlooking the cross-category or cross-domain transferable semantic features (e.g., shape, structure, etc.). Therefore, it is crucial to obtain a low-biased general annotated dataset to enhance the cross-category or cross-domain generalization capacity of the pre-trained backbones.
@@
To this end, a straight solution is to manually re-collect extensive low-biased images. However, it will not only produce expensive manual collection costs, but also inevitably incur some other undetectable bias. Recently, diffusion models [44, 43, 41] have shown powerful capacity in terms of generating high-quality synthetic images based on the text description of image contents, thus providing a feasible way to directly generate images with annotations without manual collection cost. Moreover, some studies [60, 1, 46] have demonstrated that those randomly generated images with annotations can be utilized for network training. Although most existing diffusion models can be directly utilized for general annotated dataset generation, they mainly focus on generating images with the distribution consistent with the conventional manually annotated general dataset (e.g., ImageNet) and scarcely attempt to generate low-biased images. Thus, pre-training the backbone networks on these generated general annotated datasets will not bring non-travail generalization capacity enhancement [48].
@@
To mitigate this problem, we present a low-biased general annotated dataset generation framework (lbGen), which takes the first attempt to directly generating synthetic low-biased images with category annotations. To achieve this goal, we first have to define a low-biased space where the feature representation of each image emphasizes transferable semantic characteristics. Considering that the observation that text information is closer to the ideal semantic information and the recent progress of multimodal foundation model (e.g., CLIP [40]) which aims at mapping images into a low-biased semantic space defined by language, a straight idea is to constrain the image output of the existing diffusion models to follow the semantic distribution of the specific image category in such a low-biased semantic space. Following this idea, we develop a bi-level semantic alignment loss based on the CLIP model to fine-tune the pre-trained diffusion model. In a specific, on the one hand, such a loss forces all generated images to be consistent with the semantic distribution of all categories belonging to the target dataset in the CLIP feature space using an adversarial learning scheme. On the other hand, it also requires each generated image to match the semantic description of its category name in the CLIP feature space using a simple cosine similarity metric. By doing these, we can obtain a low-biased image generation model by simply fine-tuning a pre-trained diffusion model using only all category names in the target dataset as input. In addition, to sidestep image quality degradation caused by the low-biased image generation learning, we further cast an existing image quality scoring model into a quality assurance loss to assist the bi-level semantic alignment loss for diffusion model fine-tuning.
@@
To testify the efficacy of the proposed framework, we pre-train two conventional backbone networks on our generated low-biased general annotated dataset and training specific heads on different downstream tasks. Compared with backbone networks pre-trained either on the manually collected generated dataset or that generated by existing diffusion models, our approach achieves obvious generalization performance improvement, especially when the manually annotated samples in the target task are scarce. Moreover, additional experiments prove that our pre-trained backbone networks capture lower specific bias (e.g. context, background, shape-texture), which further demonstrates the generality of our framework.
@@
  • We propose the first low-biased general annotated dataset generation framework, which jumps out of the dilemma of traditional manual data collection in terms of mitigating dataset bias.
@@
  • With our generated general low-biased dataset, the pre-trained backbone network shows state-of-the-art generalization capacity in different downstream tasks.
@@ 2.1 Datasets Bias
Since the deep learning revolution in 2012 [24], the large-scale manually collected annotated dataset (e.g., ImageNet) performs no longer a simple training dataset for its own tasks, but a general dataset utilized for backbone network pre-training which has become the indelible step for enhancing the generalization performance of various downstream tasks. However, recent studies [30, 51, 61] have consecutively revealed that these existing manually collected general dataset exhibit non-trial bias, which results in sup-optimal cross-categories and cross-domain generalization capacity, especially when the manually annotated samples in the target task are scarce. For example, Liu et al. [30] observe that deep neural networks can achieve excellent accuracy in classifying which dataset an image is from. In other words, the neural network discovers some dataset-specific patterns, a form of bias. In addition, studies in [48, 19, 37] attempt to implicitly measure the dataset bias by investigating the cross-category or cross-domain generalization capacity as well as the robustness of the models pre-trained on the dataset.
However, these works mainly focus on raising the dataset bias problem or bias measurement. In contrast, in this study, we take the first attempt to solve this problem and aims at borrowing the advantage of diffusion model in image generation to directly generate a low-biased general dataset for better backbone pre-training.
@@ 2.2 Synthetic Dataset Generation
Different from manual collection, synthetic dataset generation aims at directly generating image using deep neural networks based on some text description. For example, in early days, Zhu et al. [64, 58] utilized adversarial generative networks to model the mapping relationship between the input text description and the output image. However, these methods require large-scale high-quality images from target categories for network training, and the generalization capacity to unknown text descriptions is limited. More recently, as diffusion model shows more powerful capacity in generalization to unknown text description [43, 38, 11], some works have attempted to utilize the diffusion model to generate ImageNet-like synthetic dataset for backbone pre-training. For example, Bansal et al. [1] directly fine-tune the diffusion model on ImagNet-1K [8] and use meticulously designed prompts to generate the dataset. Lei et al. [26] utilizes ViT-GPT2 [25] to get a unique prompt to generate each image. Yuan et al. [60] resort to learn the distribution of ImageNet and use Blip-captions [27] of ImageNet as prompts to synthesize the dataset.
Although these diffusion-based methods can be utilized for general dataset generation, they mainly focus on simulating the existing ImageNet without considering the dataset bias. In this study, we attempt to fine-tune the diffusion model to directly generate a low-biased general annotated dataset without using any image-text pairs but the category names of the target dataset as input.
@@ 3 Approach
The overall training framework of lbGen is shown in Figure 2. In Section 3.1, we begin with elucidating the basic methodology employed in our training process. In Section 3.2, we then illustrate the bi-level semantic alignment module, which is the core part of our approach. Subsequently, in Section 8, we further introduce the quality assurance module for fidelity preserving the images, and finally we integrate the two components for joint learning.
@@ 3.2 Bi-Level Semantic Alignment
As mentioned in Section 1, we assume that the semantic space defined by language can be a low-biased representation and our key insight is using this characteristic of language to fine-tune a pre-trained diffusion model as our lbGen generator. We achieve this by leveraging a simple Linear-ReLU-Linear based discriminator $\mathcal{D}_{\phi}$ and utilizing CLIP to carry out a bi-level semantic alignment. We use only 1000 class names of ImageNet as our inputs, which ensures that no other biased information is introduced except semantic information of the dataset.
Entire Dataset Alignment. Building on the advantages of CLIP which has unified the image and the text into one represention space, we can initially use CLIP text encoder to extract text features $\{f_{c_{1}},f_{c_{2}},\dots,f_{c_{1000}}\}$ of classnames $\{c_{1},c_{2},\dots,c_{1000}\}$. We consider these text features to be a low-biased semantic distribution of the entire ImageNet. Then, we generate an image using prompt $c_{i}$ and send it to CLIP image encoder to get the image feature $f_{im_{i}}$. Next, we randomly choose a text feature $f_{c_{j}}$ from $\{f_{c_{1}},f_{c_{2}},\dots,f_{c_{1000}}\}$. It is worth emphasizing that we do not use the text feature which belongs to the same class as the image feature since we aim to align the whole synthetic dataset to its general semantic representation space regardless of the concrete class. Finally, features $f_{im_{i}}$ and $f_{c_{j}}$ are fed into $\mathcal{D}_{\phi}$ for computing entire semantic alignment loss as follows:
@@
Individual Image Alignment. Expect for mapping the images to be consistent with the semantic distribution of all classes within the entire dataset, we need to precisely control each category of images to match their semantic description. To this end, we introduce the individual semantic alignment loss. In particular, given the generated image ${im_{i}}$ using class name $c_{i}$, we use simple "photo of $c_{i}$" as the low-biased semantic description $p_{c_{i}}$ and send them to CLIP. Different from using CLIP to align the dataset globally in the entire semantic space, we aim at forcing the semantic information of each image to be dovetailed with its class by maximizing the cosine similarity between the image and its corresponding semantic description through CLIP. Thus, we can obtain $\mathcal{L}_{in}$ formulated as follows:
@@
By considering these two distinct levels of losses, we can finally add them together and obtain bi-level semantic alignment loss $\mathcal{L}_{bi}$ to refine the diffusion model to align more closely with the low-biased semantic reference.
@@ 3.3 Quality Assurance
Finally, we combine the losses in the bi-level semantic alignment module and quality assurance module to build up our final training object for lbGen generator as follows:
$\mathcal{L}=\mathcal{L}_{bi}+\lambda_{1}\mathcal{L}_{q},$ (8)
@@
IN-lbGen(ours) $46.3$ $62.6$ $58.0$ $71.2$ $94.1$ $86.2$ $88.6$ $68.5$ $66.0$ $74.4$
Table 1: Top-1 accuracy on transfer learning datasets. The average accuracy across eight transfer learning datasets is denoted as Avg. The best and second-best transfer learning performance of each backbone are highlighted in red and underlined. IN-SD1.5 denotes only using original SD1.5 to generate the data. We also present results on ImageNet validation set for reference.
@@ 4.1 Experimental Settings
More details about training hypersettings, data synthesis, datasets for evaluation, and computing resources are provided in Appendix.
@@ 4.2 Generalization across Downstream Tasks
Transfer Learning. Transfer learning [37] is a widely known downstream visual task and can be significantly influenced by the generalization of the pre-trained model. In our work, we aim to indicate whether the utilization of our lbGen data can enable the backbones to learn better transferable patterns. To this end, we follow fakeit [46], which uses pre-trained visual backbones as feature extractors and train simple linear logistic regression classifiers from scratch.
Figure 3: Scaling down the number of training images of eight transfer learning datasets. The benefits of using pre-trained models on our lbGen images are even more pronounced when there is less data for training.
@@
As illustrated in Table 1, we observe that models pre-trained on our dataset outperform all other candidates. Compared with the second-best synthetic datasets, we achieve $+1.4\%$ and $+3.6\%$ leading performance on ResNet50 and ViT-S respectively. More importantly, our method exhibits $1.7\%$ and $2.1\%$ average accuracy improvement and superior results in the vast majority of transfer learning datasets compared to real data. Furthermore, we investigate the transfer ability of models pre-trained on real images and our generated images when using less transfer learning training data. Such few-shot setting [52] requires an even higher generalization capacity of the model. As shown in Figure 3, it is important to note that the advantage of pre-trained models on our data is even greater when there are fewer downstream images for training. This phenomenon further underscores the lower bias of our data and the stronger generalization of the resulting model. Meanwhile, another striking finding is that achieving high accuracy on the ImageNet validation set does not necessarily correlate with enhanced cross-category generalization performance, thereby enabling us to draw more definitive conclusions regarding the dualistic impact of bias in the existing dataset on the model.
Visual Perception. Detection and segmentation are two of the most popular downstream visual perception tasks, at the same time, these two tasks can benefit a lot from a well pre-trained backbone. Hence, we want to find out if the utilization of our lbGen data can also enhance the performance on these tasks. To this end, we follow previous studies [31, 32, 47], which use Mask R-CNN [17] as detection head for COCO object detection and UperNet [57] as segmentation head for ADE20K semantic segmentation, to evaluate the performance of pre-trained ResNet50 backbone on different datasets. Moreover, to thoroughly test the generalization ability of the backbone, we progressively decrease the number of training samples and observe the outcomes.
The experimental results in Table 2 demonstrate the effectiveness of our lbGen datasets. Although pre-training on real data achieves marginally better performance with full training data, the performance of lbGen pre-trained model consistently outperforms all other pre-training data when downstream data is limited. Specifically, with only $20\%$ of the original training data, the model achieves the highest performance across both tasks, showing gains of up to $1.54\%$ APbox and $1.47\%$ mIoU compared to that pre-trained on IN-Real. This result is particularly valuable for real-world applications where collecting and annotating task-specific training data is often costly and time-consuming.